OpenStudy (anonymous):

simplify the square root of 5(8+3 the square root of 6)

OpenStudy (anonymous):

can you write it in equation form?

OpenStudy (anonymous):

I can try but my keypad doesn't have the equation symbol

OpenStudy (anonymous):

just use the equation bar beneath where you are typing

OpenStudy (anonymous):

There is no equation bar... Sqrt5(8+3sqrt6)

OpenStudy (anonymous):

like this? \[\sqrt{5}(\sqrt{8+3\sqrt{6}} )\]

OpenStudy (anonymous):

Almost just take away the square root in front of the 8

OpenStudy (anonymous):

\[\sqrt{5}\times(8+3\sqrt{6})\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Well there isn't a multiplication symbol in front of the ( but I guess tt doesnt matter

OpenStudy (anonymous):

i see.. ok, so what would you do first?

OpenStudy (anonymous):

Use distributive property

OpenStudy (anonymous):

alright, and what would that give you?

OpenStudy (anonymous):

Sqrt5(8) +sqrt5(3sqrt6)

OpenStudy (anonymous):

alright. Do you need to go into decimals? What format do you need the answer in?

OpenStudy (anonymous):

I just need to simplify it

OpenStudy (anonymous):

well, you would get